# ClearVision: Leveraging CycleGAN and SigLIP-2 for Robust All-Weather Classification in Traffic Camera Imagery

### v1

@@ V Conclusion and Future Work
In this study, we proposed a robust weather classification framework for low-quality traffic camera imagery, effectively addressing the performance degradation during nighttime conditions by combining domain adaptation through CycleGAN with Transformer-based models. Our results demonstrate that CycleGAN-enhanced domain transformations significantly improve model performance at night, with the Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + contrastive training configuration achieving the best night-time performance (85.90% accuracy) while maintaining high overall accuracy (94.00%). We showed that replacing computationally intensive CLIP with the more efficient SigLIP-2 maintains high accuracy while reducing computational demands—critical for widespread deployment.
The computational efficiency gains are particularly noteworthy, with our approach reducing training time by 89% and inference time by 83% compared to EVA-02, making it substantially more viable for resource-constrained environments. Our approach reduced the performance gap between day and night conditions from 33.81 to 8.90 percentage points in our best model, enabled by our novel weather-preserving loss in the CycleGAN framework that maintains critical visual cues during translation.
Despite these advances, limitations remain: the system struggles with extremely degraded images having near-zero illumination or severe blurring, and our relatively limited dataset constrains the models’ ability to capture rare weather patterns. Future work should explore multi-modal sensing with complementary imaging technologies, explicit temporal modeling through sequence-based architectures, expanded datasets, and edge deployment optimizations. This work addresses transportation safety challenges by providing a cost-effective solution for early, localized weather detection from existing camera infrastructure, potentially reducing weather-related traffic accidents through timely alerts. By addressing the day-night performance gap and reducing computational requirements, our approach offers a viable path toward widespread implementation of camera-based weather monitoring systems for safer transportation networks.
ACKNOWLEDGMENT
The authors would like to express their sincere gratitude to Pelmorex Weather Corporation for funding this research.

### v2

@@ V Conclusion and Future Work
In this study, we proposed a robust weather classification framework for low-quality traffic camera imagery, addressing night-time performance by combining CycleGAN with Transformer-based models. Our results show CycleGAN-enhanced transformations improve night-time performance, with Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive training achieving 85.90% night-time accuracy while maintaining 94.00% overall accuracy. By focusing contrastive loss on misclassified samples, we enhance robustness for challenging night-time images. Replacing CLIP with SigLIP-2 maintains competitive accuracy while reducing computational demands—critical for widespread deployment.
Our approach reduces training time by 89% (6 hours to 40 minutes) and inference time by 80% (15 to 3 seconds) compared to EVA-02, making it viable for resource-constrained environments. Our framework reduced the day-night performance gap from 33.81 to 8.90 percentage points, enabled by our novel weather-preserving loss in CycleGAN, which maintains critical visual cues during domain translation.
Despite these advances, limitations remain: the system struggles with images exhibiting near-zero illumination or severe blurring, and our limited dataset constrains capturing rare weather patterns. Accuracy improvements (1–2%) could be achieved through ensemble methods, combining information from multiple images with majority voting. Further enhancements could involve temporal modeling through sequence-based architectures.
Future work should explore multi-modal sensing, expanded datasets for rare conditions, and edge deployment optimization. This work addresses transportation safety by providing cost-effective early weather detection using existing camera infrastructure. Our approach offers a practical path toward widespread implementation of camera-based weather monitoring systems by narrowing the day-night performance gap, leveraging one-shot error correction, and reducing computational requirements.